# Intersections of sumsets in additive number theory

### v1

@@ Abstract.
Let AA be a subset of an additive abelian semigroup and let hAhA be the $h$-fold sumset of AA. The following question is considered: Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in the semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$. When does one have
@@ 1. The sumset intersection problem
A sequence of sets $(A_{q})_{q=1}^{\infty}$ is decreasing if $A_{q}\supseteq A_{q+1}$ for all $q\in\mathbf{N}$. A decreasing sequence of sets $(A_{q})_{q=1}^{\infty}$ is strictly decreasing if $A_{q}\neq A_{q+1}$ for all $q\in\mathbf{N}$ and asymptotically strictly decreasing if the sequence is not eventually constant, that is, if $A_{q}\neq A_{q+1}$ for infinitely many $q$.
Let AA be a subset of an additive abelian semigroup. The $h$-fold sumset of AA is the set
@@
Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets in the semigroup and let
@@
This inclusion can be proper. For example, in the semigroup $\mathbf{Z}$ of integers, if $q\in\mathbf{N}$ and
$A^{\sharp}_{q}=\{r\in\mathbf{Z}:|r|\geq q\}$
then
$n=(n+q)+(h-2)q+(-(h-1)q)\in hA^{\prime}$
@@
n=(nq)+(h2)(q)+((h1)q)-n=(-n-q)+(h-2)(-q)+((h-1)q)
for all $n\in\mathbf{N}_{0}$, and so $hA^{\sharp}_{q}=\mathbf{Z}$ for all $h\geq 2$. Let AA be any finite set of integers. The sumset hAhA is finite for all $h\in\mathbf{N}$. The asymptotically strictly decreasing sequence of sets $(A_{q})_{q=1}^{\infty}$ defined by
@@
satisfies $A=\bigcap_{q=1}^{\infty}A_{q}$. We have $hA_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$ and so
@@ 2. Intersections of sets of integers
Let AA be a set of nonnegative integers. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets of nonnegative integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$. For every positive integer $h$,
@@
We must show that if $n\in\bigcap_{q=1}^{\infty}hA_{q}$, then $n\in hA$.
The number of representations of $n$ as a sum of $h$ nonnegative integers is finite. Indeed, the set $\mathcal{R}(n)$ of $h$-tuples $(b_{1},\ldots,b_{h})\in\mathbf{N}_{0}^{h}$ such that
@@
satisfies $|\mathcal{R}(n)|=\binom{n+h-1}{h-1}$. For all $q\in\mathbf{N}$, we have $n\in hA_{q}$ and so there exists $(a_{1,q},\ldots,a_{h,q})\in A_{q}^{h}\cap\mathcal{R}(n)$ such that
@@
for infinitely many $q$ and so $b_{i}=a_{i,q}\in A_{q}$ for infinitely many $q$. Because the sequence $(A_{q})_{q=1}^{\infty}$ is decreasing, we have $b_{i}\in A_{q}$ for all $q$ and so $b_{i}\in\bigcap_{q=1}^{\infty}A_{q}=A$ and $n=b_{1}+\cdots+b_{h}\in hA$. This completes the proof. ∎
The set AA of integers is a basis of order $h$ for $\mathbf{Z}$ if $hA=\mathbf{Z}$. The set AA is a basis for $\mathbf{Z}$ if AA is a basis of order $h$ for $\mathbf{Z}$ for some positive integer $h$. If AA is a basis of order $h$ for $\mathbf{Z}$ and if $A\subseteq A^{\prime}$, then $A^{\prime}$ is a basis of order $h^{\prime}$ for $\mathbf{Z}$ for all $h^{\prime}\geq h$.
@@
Let AA be a set of integers.
  1. (1)
    If AA is a basis of order $h_{0}$ for $\mathbf{Z}$ and if $(A_{q})_{q=1}^{\infty}$ is any decreasing sequence of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$, then
@@
  2. (2)
    If AA is not a basis of order $h_{0}$ for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
@@
  3. (3)
    If AA is not a basis for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
@@
(1) Let AA be a basis of order $h_{0}$ for $\mathbf{Z}$ and let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$. For all $h\geq h_{0}$ and for all $q\in\mathbf{N}$, we have $hA=hA_{q}=\mathbf{Z}$ and so
$hA=\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
(2) If AA is not a basis of order $h_{0}$ for $\mathbf{Z}$, then $h_{0}A\neq\mathbf{Z}$. For all $q\in\mathbf{N}$, let
$A_{q}=A\cup A^{\sharp}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}.$
Then $A=\bigcap_{q=1}^{\infty}A_{q}$ and $h_{0}A_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$. It follows that
@@
(3) If AA is not a basis for $\mathbf{Z}$, then $hA\neq\mathbf{Z}$ for all $h\in\mathbf{N}$. For all $q\in\mathbf{N}$, let
$A_{q}=A\cup A^{\sharp}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}.$
Then $A=\bigcap_{q=1}^{\infty}A_{q}$ and $h_{0}A_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$. It follows that
$hA\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}$
@@
Let AA be an infinite set of integers that is bounded below and that does not contain all sufficiently large integers. There exists an asymptotically strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$.
@@
If $q\geq 1$ and $q\notin A$, then $q\in A_{q}$ but $q\notin A_{q+1}$ and so $A_{q}\neq A_{q+1}$. Because AA does not contain all sufficiently large integers, the sequence $(A_{q})_{q=1}^{\infty}$ is asymptotically strictly decreasing.
@@
Let AA be an infinite set of integers that is not a basis of order $h$ for $\mathbf{Z}$. Does there exist a strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA=\bigcap_{q=1}^{\infty}hA_{q}$?
@@
For what sets AA of integers does there exist a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
@@
For what sets AA of integers does there exist an asymptotically decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
@@ 3. Sumsets in locally compact groups
Let AA be a nonempty subset of a locally compact abelian group $G$. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of closed compact subsets of $G$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. Then
@@
Suppose that $x_{i}\notin A$ for some ii. Let $q_{0}=0$. Construct the sequence $\left(x_{i,q}\right)_{q=1}^{\infty}$ as follows;
@@
The group $G$ is locally compact. Because AA is closed and $x_{i}\notin A$, there is a compact neighborhood $K$ of $x_{i}$ that is disjoint from AA such that $x_{i,q}\in K$ for all sufficiently large $q$. Then $x_{i,q}\in K\cap A_{q}$ for all sufficiently large $q$. It follows that, for all $Q\in\mathbf{N}$, there exists $q_{1}\geq Q$ with $x_{i,q_{1}}\in K\cap A_{q_{1}}\subseteq K\cap A_{q}$ for all $q\in[1,Q]$, and so

### v2

@@ Abstract.
Let $A$ be a subset of an additive abelian semigroup and let $hA$ be the $h$-fold sumset of $A$. The following question is considered: Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in the semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$. When does one have
@@ 1. The sumset intersection problem
Let $S$ be an additive abelian semigroup and let $A$ be a subset of $S$. For every positive integer $h$, the $h$-fold sumset of $A$ is the set of all sums of $h$ not necessarily distinct elements of $A$, that is,
@@
For all $x\in S$, we denote by $r_{A,h}(x)$ the number of representations of $x$ as a sum of $h$ elements of $A$, that is,
$r_{A,h}(x)=\operatorname{\text{card}}\left\{(a_{1},\ldots,a_{h})\in A^{h}:a_{1}+\cdots+a_{h}=x\right\}.$
The subset $A$ is a basis of order $h$ for $S$ if $hA=S$ (equivalently, if $r_{A,h}(x)\geq 1$ for all $x\in S$), and a basis for $S$ if $A$ is a basis of order $h$ for $S$ for some positive integer $h$. The subset $A$ is a nonbasis of order $h$ for $S$ if $A$ is not a basis of order $h$ for $S$, that is, if $hA\neq S$. The set $A$ is a nonbasis for $S$ if $A$ is a nonbasis of order $h$ for $S$ for every positive integer $h$.
The sequence of sets $(A_{q})_{q=1}^{\infty}$ is decreasing if $A_{q}\supseteq A_{q+1}$ for all $q\in\mathbf{N}$, and strictly decreasing if $A_{q}\neq A_{q+1}$ for all $q\in\mathbf{N}$. The sequence $(A_{q})_{q=1}^{\infty}$ asymptotically strictly decreasing if the sequence is not eventually constant, that is, if $A_{q}\neq A_{q+1}$ for infinitely many $q$. Every asymptotically strictly decreasing sequence contains a strictly decreasing subsequence.
Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets in the semigroup $S$ and let
@@
This inclusion can be proper. For example, in the semigroup $\mathbf{Z}$ of integers, for every positive integer $q$, we define
$A^{\sharp}_{q}=\{r\in\mathbf{Z}:|r|\geq q\}.$
Let $h\geq 2$. For every nonnegative integer $n$, we have
$n=(n+q)+(h-2)q+(-(h-1)q)\in hA^{\sharp}_{q}$
@@
$n=(-n-q)+(h-2)(-q)+((h-1)q)\in hA^{\sharp}_{q}$
and so $hA^{\sharp}_{q}=\mathbf{Z}$.
Let $A$ be any finite set of integers. The sumset $hA$ is finite for all $h\in\mathbf{N}$ and so $hA\neq\mathbf{Z}$. The asymptotically strictly decreasing sequence of sets $(A_{q})_{q=1}^{\infty}$ defined by
@@
satisfies $A=\bigcap_{q=1}^{\infty}A_{q}$. Because $hA_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$ , we have
@@ 2. Intersections of sets of integers
Let $A$ be a set of nonnegative integers. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets of nonnegative integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$. For every positive integer $h$,
@@
We must show that for all $h\geq 2$, if $n\in\bigcap_{q=1}^{\infty}hA_{q}$, then $n\in hA$.
The number $r_{\mathbf{N}_{0},h}(n)$ of representations of $n$ as a sum of $h$ nonnegative integers is finite. Indeed, the set $\mathcal{R}(n)$ of $h$-tuples $(b_{1},\ldots,b_{h})\in\mathbf{N}_{0}^{h}$ such that
@@
satisfies $r_{\mathbf{N}_{0},h}(n)=|\mathcal{R}(n)|=\binom{n+h-1}{h-1}$. For all $q\in\mathbf{N}$, we have $n\in hA_{q}$ and so there is an $h$-tuple $(a_{1,q},\ldots,a_{h,q})\in A_{q}^{h}\cap\mathcal{R}(n)$ such that
@@
for infinitely many $q$ and so, for infinitely many $q$, we have $b_{i}=a_{i,q}\in A_{q}$ for all $i=1,\ldots,h$. Because the sequence $(A_{q})_{q=1}^{\infty}$ is decreasing, we have $b_{i}\in A_{q}$ for all $i=1,\ldots,h$ and for all $q$. It follows that $b_{i}\in\bigcap_{q=1}^{\infty}A_{q}=A$ and $n=b_{1}+\cdots+b_{h}\in hA$. This completes the proof. ∎
@@
Let $S$ be an additive abelian semigroup such that $r_{S,h}(x)$ is finite for all $h\in\mathbf{N}$ and $x\in S$. Let $A$ be a subset of $S$ and let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of subsets of $S$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. For every positive integer $h$,
$hA=\bigcap_{q=1}^{\infty}hA_{q}.$
Proof.
It suffices to observe that the proof of Theorem 1 uses only the finiteness of the representation function $R_{\mathbf{N}_{0},h}(n)$ for all $h$ and $n$. This result follows immediately. ∎
Theorem 3.
Let $S$ be an additive abelian semigroup and let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of subsets of $S$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. If $A$ is a basis of order $h_{0}$ for $S$, then
@@
Proof.
For all $h\geq h_{0}$ and for all $q\in\mathbf{N}$, we have $hA=hA_{q}=S$. and so
$hA=\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
This completes the proof. ∎
Theorem 4.
Let $A$ be a set of integers. If $A$ is a nonbasis of order $h_{0}$ for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
@@
If $A$ is a nonbasis for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
@@
For all $q\in\mathbf{N}$, let
$A_{q}=A\cup A_{q}^{\sharp}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}.$
Then $(A_{q})_{q=1}^{\infty}$ is a decreasing sequence of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$. Moreover, $h_{0}A_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$. If $A$ is a nonbasis of order $h_{0}$ for $\mathbf{Z}$, then
@@
If $A$ is a nonbasis for $\mathbf{Z}$, then for all $h\in\mathbf{N}$ we have
$hA\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
This completes the proof. ∎
Theorem 4 extends to the group of lattice points $\mathbf{Z}^{d}$ and to a more general class of semigroups. For the group $\mathbf{Z}^{d}$, we let
$A_{q}=\left\{(n_{1},\ldots,n_{d})\in\mathbf{Z}^{d}:|n_{i}|\geq q\text{ for all }q\in\mathbf{N}\right\}.$
Theorem 5.
Let $S$ be an additive abelian semigroup such that $S$ contains a strictly decreasing sequence of sets $\left(A_{q}^{\sharp}\right)_{q=1}^{\infty}$ such that $\bigcap_{q=1}^{\infty}A_{q}^{\sharp}=\emptyset$ and $hA_{q}^{\sharp}=S$ for all $h\geq 2$ and $q\geq 1$.
If $A$ is a nonbasis of order $h_{0}$ for $S$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
$h_{0}A\neq\bigcap_{q=1}^{\infty}hA_{q}.$
If $A$ is a nonbasis for $S$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
$hA\neq\bigcap_{q=1}^{\infty}hA_{q}$
for all $h\in S$.
Proof.
If $A$ is a nonbasis of order $h_{0}$ for $S$, then $h_{0}A\neq S$. For all $q\in\mathbf{N}$, let
$A_{q}=A\cup A_{q}^{\sharp}.$
Then $A=\bigcap_{q=1}^{\infty}A_{q}$ and $h_{0}A_{q}=S$ for all $q\in\mathbf{N}$. It follows that
$h_{0}A\neq S=\bigcap_{q=1}^{\infty}h_{0}A_{q}.$
If $A$ is a nonbasis for $S$, then for all $h\in\mathbf{N}$ we have $hA\neq S$ and
$hA\neq S=\bigcap_{q=1}^{\infty}hA_{q}$
@@
Let $A$ be an infinite set of integers that is bounded below and that does not contain all sufficiently large integers. There exists an asymptotically strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$.
@@
If $q\geq 1$ and $q\notin A$, then $q\in A_{q}$ but $q\notin A_{q+1}$ and so $A_{q}\neq A_{q+1}$. Because $A$ does not contain all sufficiently large integers, the sequence $(A_{q})_{q=1}^{\infty}$ is asymptotically strictly decreasing.
@@
Let $A$ be an infinite set of integers that is a nonbasis of order $h$ for $\mathbf{Z}$. Does there exist a strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA=\bigcap_{q=1}^{\infty}hA_{q}$?
@@
For what sets $A$ of integers does there exist a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
@@
For what sets $A$ of integers does there exist an asymptotically decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
@@ 3. Sumsets in locally compact groups
Let $A$ be a nonempty subset of a locally compact abelian group $G$. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of closed compact subsets of $G$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. Then
@@
Suppose that $x_{i}\notin A$ for some $i$. Let $q_{0}=0$. Construct the sequence $\left(x_{i,q}\right)_{q=1}^{\infty}$ as follows;
@@
The group $G$ is locally compact. Because $A$ is closed and $x_{i}\notin A$, there is a compact neighborhood $K$ of $x_{i}$ that is disjoint from $A$ such that $x_{i,q}\in K$ for all sufficiently large $q$. Then $x_{i,q}\in K\cap A_{q}$ for all sufficiently large $q$. It follows that, for all $Q\in\mathbf{N}$, there exists $q_{1}\geq Q$ with $x_{i,q_{1}}\in K\cap A_{q_{1}}\subseteq K\cap A_{q}$ for all $q\in[1,Q]$, and so
@@ 4. Semigroup problems
I thank Diego Marques for the observation that led to Theorem 2.